find the inverse of the following matrix A and rank of a 2. 1. 2

1. 3. 0

-1. 1. 2


1
Expert's answer
2022-01-30T15:16:09-0500

Augmenting A with a "3\u00d73" identity matrix


"=\begin{pmatrix}
 2 & 1&2| &1&0&0\\
 1&3&0| & 0&1&0\\
-1&1&2|&0&0&1
\end{pmatrix}"


"\frac{1}{2}R_1\to\>R_1"

"-R_3\to\>R_3"


"\begin{pmatrix}
 1&\frac{1}{2}&1&|& \frac{1}{2}&0&0 \\
 1&3&0&|& 0&1&0\\
1&-1&-2&|&0&0&-1
\end{pmatrix}"


"R_2-R_1\to\>R_2"

"R_3-R_1\to\>R_3"



"\begin{pmatrix}
 1&\frac{1}{2}&1&|&\frac{1}{2}&0&0 \\
 0&\frac{5}{2}&-1&|&\frac{-1}{2}&1&0 \\
0&\frac{-3}{2}&-3&|&\frac{-1}{2}&0&-1
\end{pmatrix}"


"\frac{2}{5}R_2\to\>R_2"

"\frac{-2}{3}R_3\to\>R_3"



"\begin{pmatrix}
 1&\frac{1}{2}&1&|&\frac{1}{2}&0&0 \\
 0&1&\frac{-2}{5}&|&\frac{-1}{5}&\frac{2}{5}&0 \\
0&1&2&|&\frac{1}{3}&0&\frac{2}{3}
\end{pmatrix}"


"\frac{5}{12}(R_3-R_2)\to\>R_3"

"R_2+\frac{2}{5}R_3\to\>R_2"


"\begin{pmatrix}
 1&\frac{1}{2}&1&|&\frac{1}{2}&0&0\\
 0&1&0&|&\frac{-1}{9} & \frac{1}{3}&\frac{1}{9}\\
0&0&1&|&\frac{2}{9}&\frac{-1}{6}&\frac{5}{18}
\end{pmatrix}"



"R_1-R_3\to\>R_3"


"\begin{pmatrix}
 1&\frac{1}{2}&0&|&\frac{5}{18}&\frac{1}{6}&\frac{-5}{18} \\
 0&1&0&|&\frac{-1}{9} & \frac{1}{3}&\frac{1}{9}\\
0&0&1&|&\frac{2}{9}&\frac{-1}{6}&\frac{5}{18}
\end{pmatrix}"



"R_1-\frac{1}{2}R_2\to\>R_1"


"=\begin{pmatrix}
 1&0&0| & \frac{1}{3}&0&\frac{-1}{3}\\
 0&1&0| & \frac{-1}{9}&\frac{1}{3}&\frac{1}{9}\\
0&0&1|&\frac{2}{9}&\frac{-1}{6}&\frac{5}{18}
\end{pmatrix}"


Inverse is matrix on the right of augmented matrix


"\therefore\>A^{-1}=\begin{pmatrix}
 \frac{1}{3}& 0&\frac{-1}{3} \\
 \frac{-1}{9}&\frac{1}{3} & \frac{1}{9}\\
\frac{2}{9}&\frac{-1}{6}&\frac{5}{18}
\end{pmatrix}"



Rank is number of non-zeros row on the left of the augmented matrix


"\therefore" Rank "=3"